# QSHS: An Axion Dark Matter Resonant Search Apparatus

### v1

@@ Abstract
[…]
@@ 1 Physics Motivation
×$\nabla\times({\bf B}/\mu_{0})-\partial_{t}(\varepsilon_{0}{\bf E})={\bf j}_{e}+{\bf j}_{a},$ (3)
@@
×$\times\left(\frac{\rho_{a}}{\rm 0.45\,GeV/cc}\right)\left(\frac{\nu_{a}}{\rm 750\,MHz}\right)\left(\frac{Q}{70,000}\right),$ (8)
@@
[…]
@@
×$\times\left(\frac{\rho_{a}}{\rm 0.45\,GeV/cc}\right)\left(\frac{\nu_{a}}{\rm 5\,GHz}\right)\left(\frac{Q}{3,000}\right).$ (10)
@@ 3 Cryostat
[…]
@@ 5 Magnet
The magnet was supplied integrated into the cryostat by Oxford Instruments. It is conductively cooled via the copper thermal shield connecting the outer diameter of its top flange to the PT2 plate. The magnet is a dry superconducting design. The inner coil is niobium-tin, and the outer two coils are niobium-titanium. Current is supplied to the three coils in series. A heat switch permits persistent operation. External power supplies provided by Oxford intruments provide the $\rm 160\,A$ of current necessary to reach full field, which is $\rm 8\,T$ at the centre of the bore. The field is nominally cylindrically symmetric, although outside the magnet bore the off-centre quantum electronics field shield breaks the cylindrical symmetry in its vicinity. Figure 4 is a simulated field map inside the magnet provided by Oxford Instruments.
The stray field is kept minimal by the outer niobium-tin windings, which act as a magnetic field shield. The stray field falls to $\rm 5\,G$ at $\rm 1.2\,m$ and $\rm 1.4\,m$, respectively, in the horizontal plane and along the axis from the centre of the magnet bore. The well confined high field region ensures that stray field does not affect neighbouring labs and corridors, or couple to the rebar in the concrete floor. It also provides a low field environment for workers and for electronics connected the fridge, allowing cable lengths to be minimal, and permitting operators to work without concerns about field over the majority of the lab and at all the control racks. Short cables are advantageous for minimising microwave signal losses and for preserving the sawtooth waveforms used to drive the piezo motors described in Section 6.2.
@@ 6.1 The QSHS Cavity
[…]
@@ 7 Receiver Electronics
Figure 7 is a schematic of the receiver electronics connected to the resonant cavity system. The variable-coupling port to the cavity feeds an ultra-low-noise amplifier housed in the magnetic field shield via a Low Noise Factory [42] LNF-CIC4_8A circulator. The ultra-low-noise amplifier may be one of several under development in the group, or this stage may be bypassed, in which case the signal feeds directly to the input of the first cryogenic HFET amplifier mounted at the $\rm 4\,K$ stage. A second Low Noise Factory[42] LNF-CIC4_8A circulator provides isolation between the ultra-low-noise amplifier output and the input to the HFET amplifier stages. The HFET amplification consists of two Low Noise Factory LNF-LNC4_8G [42] amplifiers in series, each having a bandwidth of $\rm 4-8\,GHz$ and providing $\rm 40\,dB$ of gain. The output of the second of these HFET amplifiers is fed through a vacuum feed-through to a Low Noise Factory [42] LNF-LNR4_8F room temperature microwave amplification stage, providing a further $\rm 44\,dB$ of gain at room temperature. A cryogenic RF switch allows for swept transmission measurements of the cavity between the weak and variable ports or a reflection measurement off the variable port to verify critical coupling. A room temperature RF switch allows the receiver chain output to feed either the second port of an Anritsu vector network analyser or the room temperature heterodyne receiver apparatus. A Polyphase Microwave [43] IRM4080B image-reject mixer down-converts power from a band around the cavity $\rm TM_{010}$ mode frequency $\nu_{0}$ to an intermediate frequency (IF) centered at $\rm 10.7\,MHz$.
@@ 7.1 Field Shield
Figure 8: The QSHS field shield. (a) Rendered cross section of the shield, with layers marked. Indicated dimensions are in mm. Inset shows a partly disassembled view. (b) Shield in position beneath the mixing chamber, with superimposed stray field contours (simulated without the shield). Labels indicate field magnitude in mT.
@@
The shield is cooled with the primary magnet unenergized and therefore experiences the Earth’s magnetic field $B_{\mathrm{earth}}\approx 50\,\mu$T. Most of this field is routed through the outer ferromagnetic layer, ensuring that when the outer superconducting layer is cooled through its transition, the field it encloses is much smaller than $B_{\mathrm{earth}}$. To suppress vortex trapping in the superconductor, its cooling path runs through a piece of brass, whose comparatively low thermal conductance should slow down the superconducting transition. After the cryostat reaches base temperature, the primary coil is ramped up. By this time, the outer superconducting layer is well below its transition temperature, so its $B_{c}$ is much larger than the stray field it experiences. Any flux that does penetrate it is primarily routed through the inner ferromagnetic layer rather than reaching the amplifiers. The final passive shielding comes from the inner superconducting layer, which does not reduce the field magnitude but does impose a boundary condition that it runs parallel to its surface and therefore predominantly along the cylinder axis. If this condition is achieved, then the requirement that the $\mathbf{B}$-field have zero divergence and curl means that it will be uniform in this space.
@@
When the primary coil is ramped up, the outer ferromagnetic layer will saturate and therefore lose some shielding effectiveness, but the superconducting layers should remain below their transition fields. Only a small part of the stray field should therefore penetrate the outer superconducting and inner ferromagnetic shields to reach the amplifiers when they are operating.
@@
$\mathbf{F}$ $=\gradient(\mathbf{m}\cdot\mathbf{B}_{\mathrm{stray}})$
$=\frac{8\pi}{\mu_{0}}R^{3}B_{\mathrm{stray}}\gradient B_{\mathrm{stray}}.$ (15)
At the location of the shield, the stray field of the magnet is $B_{\mathrm{stray}}\approx 10^{-2}~{}\mathrm{T}$ with $\left|\gradient B_{\mathrm{stray}}\right|\approx 10^{-1}~{}\mathrm{T}~{}\mathrm{m}^{-1}$ and so, taking $R=0.07~{}\mathrm{m}$ from Figure 8(a), the magnitude of the force is
@@ 8 Data Acquisition
With the tuning rod in a fixed position and the variable coupling antenna inserted such that the electronics is critically coupled to the cavity $\rm TM_{010}$ mode, data acquisition consists, first, of wide and narrow band sweeps of the cavity in transmission to determine mode resonant frequency and quality factor. Second, data is acquired through the heterodyne receiver described in Section 7. The raw data consists of Welch estimates of the power spectrum over the bandwidth of the cavity mode. Having acquired sufficient data to achieve a signal to noise ratio, as defined in Equation (11), of $\rm SNR=4$, for the signal model under examination, the cavity transmission functions, power spectra, and slow controls data are stored in hdf5 format for off-line analysis. The piezo motors are then used to move the tuning rod, and the cycle repeats until the desired range of mode frequencies and axion masses has been probed.
@@ 9 Analysis
10 Summary
In this paper we have described the initial hardware configuration of the QSHS resonant cavity axion search. The instrument is designed both as a test bed for new low-noise quantum instrumentation and as a search for pseudoscalar hidden-sector dark matter at sensitivities comparable with the expectations for QCD axions. The base temperature of the cavity is $\rm 18.5\,mK$ so that, instrumented with quantum electronics, the ratio of the thermal equilibrium energy, $k_{B}T$, to the energy of a single quantum of excitation of the resonator, $h\nu$, should be less than unity. The apparatus has room for expansion for other experiments and tests of novel experimental techniques. Tests of quantum electronics supplied to the QSHS collaboration and searches for axions using the apparatus will be the subject of future collaboration publications.
@@ 11 Acknowledgements
The QSHS collaboration is funded by the Science and Technology Facilities Council (STFC), Quantum Technologies for Fundamental Physics (QTFP) programme, grant codes ST/T006811/1 (Sheffield), ST/T006064/1 (NPL), ST/T006099/1 (UCL), ST/T006102/1 (Lancaster), ST/T006145/1 (Liverpool, Oxford), ST/T006277/1 (Oxford), ST/T006625/1,2 (Oxford, Cambridge) and ST/T006242/1 (RHUL). Further support was provided to the associated ‘ParaPara’ project, funded by STFC, again as part of the QTFP programme, grant codes ST/W006502/1 (Lancaster) and ST/W006464/1 (UCL). We wish to thank the Mary and Arthur Hogg Scholarship and Sheffield Alumni Fund for financial support for A. Sundararajan.

### v2

@@ Abstract
We describe a resonant cavity search apparatus for axion dark matter constructed by the Quantum Sensors for the Hidden Sector (QSHS) collaboration. The apparatus is configured to search for QCD axion dark matter, though also has the capability to detect axion-like particles (ALPs), dark photons, and some other forms of wave-like dark matter. Initially, a tuneable cylindrical oxygen-free copper cavity is read out using a low noise microwave amplifier feeding a heterodyne receiver. The cavity is housed in a dilution refrigerator and threaded by a solenoidal magnetic field, nominally $\rm 8\,T$. The apparatus also houses a magnetic field shield for housing superconducting electronics, and several other fixed-frequency resonators for use in testing and commissioning various prototype quantum electronic devices sensitive at a range of axion masses in the range $\rm 2.0$ to $\rm 40\,{\mathrm{\mu}}eV/c^{2}$. The apparatus as currently configured is intended as a test stand for electronics over the relatively wide frequency band attainable with the $\rm TM_{010}$ cavity mode used for axion searches. We present performance data for the resonator, dilution refrigerator, and magnet, and plans for the first science run.
@@ 1 Physics Motivation
$\nabla\times({\bf B}/\mu_{0})-\partial_{t}(\varepsilon_{0}{\bf E})={\bf j}_{e}+{\bf j}_{a},$ (3)
@@
$\times\left(\frac{\rho_{a}}{\rm 0.45\,GeV/cc}\right)\left(\frac{\nu_{a}}{\rm 750\,MHz}\right)\left(\frac{Q}{70,000}\right),$ (8)
@@
$f_{\rm nlm}=\frac{\left(\int{\bf E}_{\rm nlm}(x)\cdot{\bf B}(x)\,dV\right)^{2}}{\left(\int|{\bf E_{\rm nlm}}|^{2}\,dV\right)\left(\int|{\bf B}|^{2}\,dV\right)},$ (9)
@@
$\times\left(\frac{\rho_{a}}{\rm 0.45\,GeV/cc}\right)\left(\frac{\nu_{a}}{\rm 5\,GHz}\right)\left(\frac{Q}{3,000}\right).$ (10)
@@ 3 Cryostat
The dilution refrigerator requires an 18 litre charge of $\rm{}^{3}He$ in a $\rm 100$ litre vessel of $\rm{}^{3}He/^{4}He$ gas mixture supplied by Oxford Instruments. Figure 1 shows the positions of the still, cold, and mixing chamber plates. The cooling power of the fridge at the mixing chamber plate is $\rm 450\,\mu W$ at $\rm 100\,mK$ and $\rm 12\,\mu W$ at $\rm 20\,mK$. A Lakeshore 372 AC resistance bridge and Lakeshore ruthenium oxide temperature sensors allow for cryogenic temperature control [31].
@@ 5 Magnet
The magnet was supplied integrated into the cryostat by Oxford Instruments. It is conductively cooled via the copper thermal shield connecting the outer diameter of its top flange to the PT2 plate. The magnet is a dry superconducting design. The inner coil is niobium-tin, and the outer two coils are niobium-titanium. Current is supplied to the three coils in series. A heat switch permits persistent operation. External power supplies provided by Oxford intruments provide the $\rm 160\,A$ of current necessary to reach full field, which is $\rm 8\,T$ at the centre of the bore. The field is nominally cylindrically symmetric, although outside the magnet bore the off-centre quantum electronics field shield breaks the cylindrical symmetry in its vicinity. Figure 4 is a simulated field map inside the magnet provided by Oxford Instruments. The fields of superconducting magnets in persistent mode are very time-stable, and we do not anticipate contributions to the noise from fluctuations in the magnetic field itself. Care must however be taken to prevent wires from moving in the applied field. In particular, the low frequency cables connected to temperature sensors can pick up noise due to induced currents. Such cables are therefore tied firmly to supports and fixtures to minimise such vibrations.
The stray field is kept minimal by the outer niobium-titanium windings, which act as a magnetic field shield. The stray field falls to $\rm 5\,G$ at $\rm 1.2\,m$ and $\rm 1.4\,m$, respectively, in the horizontal plane and along the axis from the centre of the magnet bore. The well confined high field region ensures that stray field does not affect neighbouring labs and corridors, or couple to the rebar in the concrete floor. It also provides a low field environment for workers and for electronics connected the fridge, allowing cable lengths to be minimal, and permitting operators to work without concerns about field over the majority of the lab and at all the control racks. Short cables are advantageous for minimising microwave signal losses and for preserving the sawtooth waveforms used to drive the piezo motors described in Section 6.2.
@@ 6.1 The QSHS Cavity
Cavity simulations herein were conducted using ANSYS electromagnetic finite element analysis simulations, primarily ANSYS HFSS [32]. The QSHS cavity has a tuneable range 4.1 to 7.5 GHz, corresponding to an axion search range approximately $\rm 17$ to $\rm 31\,\mu eV/c^{2}$. This range was chosen to optimise the operating frequencies to match with devices being developed by the wider QSHS group. The copper cavity mount also has space for three static frequency cavities for further device testing and development. Simulation results and overlaid room temperature measurements of the cavity $\rm TM_{010}$ mode frequency obtained in transmission using a network analyser are shown in Figure 6.
@@ 7 Receiver Electronics
Figure 7 is a schematic of the receiver electronics connected to the resonant cavity system. Vendor web sites supplied in the references contain detailed information of each component in the receiver chain. The variable-coupling port to the cavity feeds an ultra-low-noise amplifier housed in the magnetic field shield via a Low Noise Factory [42] LNF-CIC4_8A circulator. The ultra-low-noise amplifier may be one of several under development in the group, or this stage may be bypassed, in which case the signal feeds directly to the input of the first cryogenic HFET amplifier mounted at the $\rm 4\,K$ stage. A second Low Noise Factory [42] LNF-CIC4_8A circulator provides isolation between the ultra-low-noise amplifier output and the input to the HFET amplifier stages. The HFET amplification consists of two Low Noise Factory LNF-LNC4_8G [42] amplifiers in series, each having a bandwidth of $\rm 4-8\,GHz$ and providing $\rm 40\,dB$ of gain. The output of the second of these HFET amplifiers is fed through a vacuum feed-through to a Low Noise Factory [42] LNF-LNR4_8F room temperature microwave amplification stage, providing a further $\rm 44\,dB$ of gain at room temperature. A cryogenic RF switch allows for swept transmission measurements of the cavity between the weak and variable ports or a reflection measurement off the variable port to verify critical coupling. A room temperature RF switch allows the receiver chain output to feed either the second port of an Anritsu vector network analyser or the room temperature heterodyne receiver apparatus. A Polyphase Microwave [43] IRM4080B image-reject mixer down-converts power from a band around the cavity $\rm TM_{010}$ mode frequency $\nu_{0}$ to an intermediate frequency (IF) centered at $\rm 10.7\,MHz$.
The dominant sources of noise in cavity axion searches are Johnson noise from the cavity walls and broadband electronic noise added by the components of the receiver electronics. We therefore aspire to run as cold of a resonator as we can, and to develop and use the lowest possible noise readout electronics. The noise contribution of each component in the receiver chain is divided by the gain of all the previous amplification stages, so that in practice only the first couple of amplification stages make significant contributions to the overall noise temperature.
Calibration of the apparatus consists of conversion of power levels to physical temperatures. Primary calibration will consist of a measurement of the noise temperature of the first Low Noise Factory HFET amplifier. The hot-cold-load method will be used. Having measured the noise temperature of this amplifier, we can switch in an upstream ultra-low-noise device under test (DUT). We employ a technique that has become common in the analysis of receiver electronics. We inject a narrow signal in the centre of the receiver band, through a 50 ohm fixed attenuator to which is afixed a temperature sensor and a heater. We measure the ratio of the power of this peak to the noise floor in the surrounding band as a function of temperature, first with and then without the DUT connected between the attenuator and the first HFET amplifier. The amount by which this ratio increases is the so-called ’signal-to-noise ratio improvement’ (SNRI) [44]. The SNRI measurement is then used to infer by how much the first low noise stage is improving the noise temperature of the receiver chain. In reality, it will be a multi-element equivalent circuit whose parameters are constrained by the measured SNRI. Details of the calibration procedure and results are topics for discussion in future publications.
@@ 7.1 Field Shield
Figure 8: The QSHS field shield. (a) Rendered cross section of the shield, with layers marked. Indicated dimensions are in $m$. Inset shows a partly disassembled view. (b) Shield in position beneath the mixing chamber, with superimposed stray field contours (simulated without the shield). Labels indicate field magnitude in mT.
@@
[…]
@@
When the primary coil is ramped up, the outer ferromagnetic layer will saturate and therefore lose some shielding effectiveness, but the superconducting layers should remain below their transition fields. Only a small part of the stray field should therefore penetrate the outer superconducting and inner ferromagnetic shields to reach the amplifiers when they are operating. If it turns out that amplifiers under test do not work in the shield, it is an option to obtain a magnetic field probe and measure the field inside the shield as a diagnostic.
@@
$\mathbf{F}$ $=\vec{\nabla}(\mathbf{m}\cdot\mathbf{B}_{\mathrm{stray}})$
$=\frac{8\pi}{\mu_{0}}R^{3}B_{\mathrm{stray}}\vec{\nabla}B_{\mathrm{stray}}.$ (15)
At the location of the shield, the stray field of the magnet is $B_{\mathrm{stray}}\approx 10^{-2}~\mathrm{T}$ with $\left|\vec{\nabla}B_{\mathrm{stray}}\right|\approx 10^{-1}~\mathrm{T}~\mathrm{m}^{-1}$ and so, taking $R=0.07~\mathrm{m}$ from Figure 8(a), the magnitude of the force is
@@ 8 Data Acquisition
With the tuning rod in a fixed position and the variable coupling antenna inserted such that the electronics is critically coupled to the cavity $\rm TM_{010}$ mode, data acquisition consists, first, of wide and narrow band sweeps of the cavity in transmission to determine mode resonant frequency and quality factor. Second, data is acquired through the heterodyne receiver described in Section 7. The raw data consists of Welch estimates of the power spectrum over the bandwidth of the cavity mode. Having acquired sufficient data to achieve a signal to noise ratio, as defined in Equation (11), of $\rm SNR=4$, for the signal model under examination, the cavity transmission functions, power spectra, and slow controls data are stored in hdf5 format for off-line analysis. The piezo motors are then used to move the tuning rod, and the cycle repeats until the desired range of mode frequencies and axion masses has been probed. A good target is to aim for 90% data taking live time, opening the apparatus approximately every three months for inspection and diagnostics. If drifts are observed in mode characteristics during long integration times, these times will be broken up into segments and additional measurements of the cavity mode characteristics will be added.
@@ 9 Analysis
10 Summary and Next Steps
In this paper we have described the initial hardware configuration of the QSHS resonant cavity axion search. The instrument is designed both as a test bed for new low-noise quantum instrumentation and as a search for pseudoscalar hidden-sector dark matter at sensitivities comparable with the expectations for QCD axions. The base temperature of the cavity is $\rm 18.5\,mK$ so that, instrumented with quantum electronics, the ratio of the thermal equilibrium energy, $k_{B}T$, to the energy of a single quantum of excitation of the resonator, $h\nu$, should be less than unity.
Next steps include calibration of the instrument sensitivity through measurement of the noise temperature of the receiver. This will involve measuring the noise temperature of the chain starting at the first HFET amplifier. This noise temperature should be dominated by the noise temperature of the first HFET stage. This result can be used to assess the change in noise temperature when additional device prototypes are added before this first HFET amplifier. A low measured noise temperature is an indication that the gain propagated through the receiver chain is sufficient to maintain adequate signal-to-noise ratio at the point where the measured signal is digitised. Such a measurement also provides an absolute calibration of the receiver chain, so that any power excess can be converted to an absolute signal power deposited in the cavity. The apparatus has room for expansion for other experiments and tests of novel experimental techniques. Future experimental configurations will optimise the choice of resonator to optimise signal sensitivity, matched to the receiver electronics. These and other tests of quantum electronics supplied to the QSHS collaboration and searches for axions using the apparatus will be the subject of future collaboration publications.
@@ 11 Acknowledgements
The QSHS collaboration is funded by the Science and Technology Facilities Council (STFC), Quantum Technologies for Fundamental Physics (QTFP) programme, grant codes ST/T006811/1 (Sheffield), ST/T006064/1 (NPL), ST/T006099/1 (UCL), ST/T006102/1 (Lancaster), ST/T006145/1 (Liverpool, Oxford), ST/T006277/1 (Oxford), ST/T006625/1,2 (Oxford, Cambridge) and ST/T006242/1 (RHUL). Further support was provided to the associated ‘ParaPara’ project, funded by STFC, again as part of the QTFP programme, grant codes ST/W006502/1 (Lancaster) and ST/W006464/1 (UCL). The LLNL group is supported by the U.S. Department of Energy under Contract DE-AC52-07NA27344. LLNL-JRNL-2007312. We wish to thank the Mary and Arthur Hogg Scholarship and Sheffield Alumni Fund for financial support for A. Sundararajan.